# Catch Me if You Can: Detecting Unauthorized Data Use in Deep Learning Models

### v1

@@ 6 Evaluation on ImageNet Training
MembershipTracker can still maintain 0% FPR@100% TPR, when the marking ratio is reduced from 0.01% to 0.005%. However, when the marking ratio is further reduced to 0.0025%, MembershipTracker experiences lower auditing performance - this trend is similar to that in our ablation study in Appendix C.1.1 and thus we defer the detailed discussion to Appendix C.1.1 (due to space constraints). Nevertheless, even under such a challenging case, MembershipTracker still achieves reasonably good performance as shown in the blue line of Fig. 10; we leave the improvements to future study.
Figure 10: MembershipTracker under full-sized ImageNet training with different portions of samples for data marking. Aggressively reducing the portion of samples marked by each user can lead to reduced performance, though MembershipTracker can still maintain reasonably high auditing performance even when each user can only mark merely 0.0025% of the training data (the blue line).
7 Limitations
There are three main limitations by MembershipTracker, and we discuss them next.
@@ 7 Limitations
In order to trace data provenance in a given model, we assume the model is trained on a small number of target samples that are marked with MembershipTracker. This is important in ensuring that the target samples can be strongly memorized by the model, and the set-based MI process can leverage the collective information across the target samples to generate a reliable outcome.
However, this assumption may be invalid if: (a) the users have limited samples available for data marking; or (b) they have marked several samples, but only a subset of them are used for training the model. We discuss both cases next.
(a) The first case is dependent on the application scenario and may be overcome by the users themselves. E.g., if the users want to know if their images posted on social media platform may be misused, they can upload more images that are marked with MembershipTracker (both Section 6 and Appendix C.1.1 study how different number of samples marked by the users may affect MembershipTracker’s performance).
With that said, marking more samples does not mean all of them would be used for training the model, which leads to the second threat to the validity of MembershipTracker.
(b) Recall the set-based MI process is underpinned by the observation that the average loss of the samples by the non-target users are higher than that of the target users (Section 4.4.2). This, however, might not hold, if only a subset of the marked samples from each target user are used for model training. This is because those samples that are not used for training would yield higher prediction loss, which in turn increases the average loss on the entire set of marked samples, and affects MembershipTracker.
Table 5: Evaluation on the partial inclusion of the marked samples in model training. With lesser samples (50 samples per user) included for training, the average loss on the entire set of marked samples by each target user increases, which in turn increases the FPR by MembershipTracker.
% of marked samples FPR@100% TPR FPR@100% TPR
used for training (100 target users) (5 target users)
@@
90% (45/50) 0.02 0.00
70% (35/50) 0.02 0.04
50% (25/50) 0.28 0.34
30% (15/50) 20.12 2.46
20% (10/50) 76.34 6.82
Evaluation. To understand this, we perform an experiment where we assume the model owner collects 50 samples from each user (0.2% of the training set), and he/she deliberately includes only a subset of samples from each target user for training. Table 5 shows the results, where we observe a similar trend when MembershipTracker is applied to support 100 target users (a more challenging setting) or 5 target users only. In both cases, the FPR incurred by MembershipTracker grows as the number of marked samples used for training decreases, which is due to the higher average loss on the target users’ samples as we explained earlier.
Overall, we find that MembershipTracker can still maintain $<0.5\%$ FPR@100% TPR, even when only 50% of the marked samples by each target user are used for training. But its performance drops rapidly when lesser samples are included for training. Nevertheless, we remark that using only a subset of the user data for training merely represents a special usage case for understanding the potential threat to MembershipTracker’s performance, and it does not necessarily represent the real-world practice of model training.
Specifically, given that DL models are most effective when applied to large datasets [60, 5], it is often in the model creator’s interest to use more data (e.g., use all the collected data) for training, and MembershipTracker can achieve superior performance in such a setting. As the first of its type, how MembershipTracker may reshape the current model training practice remains to be seen.
@@
Secondly, despite our best effort in keeping the targeted changes from severely distorting the marked data, the artifacts created by MembershipTracker are still visible under close inspection (more visual examples are shown in Appendix A). However, performing manual inspection to filter out the specially-marked samples by MembershipTracker can still be a challenge, particularly under the large-scale training setting (e.g., with millions of images). With that said, creating more subtle targeted changes still remains an important venue for future investigation.
@@
Finally, since MembershipTracker targets the model’s propensity in memorizing data, in principle, any effective measures that can mitigate the model’s memorization on training data would be able to degrade the auditing performance by MembershipTracker. Indeed, our evaluation in Section 5.2 shows that common approaches such as differential privacy and data augmentation can be configured to degrade the auditing success by MembershipTracker, but they also result in utility loss to the model (undesirable).
@@
8 Discussion and Future Work
This work presents MembershipTracker, a practical data provenance tool that can support ordinary users to audit whether their specific data are used to train deep learning models without their permission. If their data are found to be misused, the users can take legal action or request the “right to be forgotten” in accordance with privacy regulations such as GDPR [2] (e.g., based on techniques on machine unlearning [7]). Finally, we discuss two other future work directions.
First, while MembershipTracker retrofits the conventional membership inference (attack) to enable responsible AI, it also faces the risk of being misused by the malicious party. Notably, the data marked with MembershipTracker would now become easier to be de-identified by any party with access to them. However, MembershipTracker should be applied only when the users do not want their data to be used for training DL models, in which membership privacy may not be of the users’ concern. By contrast, if the users are willingly sharing their data for model training, they should refrain from using MembershipTracker to avoid unnecessary membership exposure. Future studies can explore potential usage cases where MembershipTracker is applicable and protecting the membership privacy of the marked samples is necessary, and consider the further adaptation of MembershipTracker to cater to such a need.
Next, future work can explore the potential of extending MembershipTracker to other domains like generative models. Prior work show that performing MI on generative image models (e.g., Stable Diffusion) is also related to the model’s propensity in memorizing individual data points [9]. Thus, while MembershipTracker demonstrates promising performance under the million-scale training setting, future study can investigate whether the large generative models would also be prone to memorize the samples marked with MembershipTracker under the billion-scale training environment.
@@ Appendix A Model Training Details
For CIFAR10 and CIFAR100, we train a WideResNet-28-4 model on 25,000 samples for 100 epochs, with a learning rate of 0.1 (decayed by 5 at epoch 60, 80, 90), a weight decay of 5e-4 and momentum of 0.9. For TinyImageNet and ArtBench, we fine-tune an ImageNet-pretrained ResNet-18 model on 25,000 samples. For CelebA, we first follow [1] to create a 307-class subset from the dataset, and then similarly fine-tune an ImageNet-pretrained model on 2,000 samples. For fine-tuning, we use a small fine-tuning rate of 0.01 with 30 epochs and a momentum of 0.9. In all datasets except CelebA (where we find that fine-tuning on the entire set yields better accuracy), we use 20% of the training data as the validation set. We also use common data augmentations such as random cropping and horizontal flip.
@@
Table 6: Evaluation on different models (parenthesized numbers show accuracy diff. compared with the original models).
@@
Table 7: Evaluation on different training-set sizes.
@@ Appendix C Ablation Study
MembershipTracker consists of a data marking and set-based MI process, and we conduct a detailed ablation study into these two in Appendix C.1 and C.2.
C.1 Two-step Data Marking
@@
Table 8: Ablation study of MembershipTracker’s two-step data marking process: 1ga1g: Image blending with OOD feature; 1gb1g: Perlin noise injection. Each column reports the performance under a specific number of target samples contributed by the users (5$\sim$15 samples by each).
@@
(2) From left to the right: MembershipTracker’s auditing performance degrades as the number of samples contributed by the users reduces. However, even in the challenging case where each user marks only 5 samples, the full MembershipTracker still incurs only 0.64% FPR@100% TPR.
@@
C.1.1 Investigating the two-step data marking process
@@
We consider a total of 100 target users, and each user contributes different number of samples (5 to 15) to the training set (25,000 instances). Table 8 reports the results, and there are two main findings.
@@
Meanwhile, the performance difference between different approaches in Table 8 shrinks as the number of target samples contributed by the users increases. Thus, there is a low FPR even when auditing the original target samples (without data marking), e.g., 0.34% FPR@100% TPR on the setting with 15 target samples per user. This is in fact due to the effectiveness of the proposed set-based MI process (without which, performing the standard instance-based MI incurs 44% FPR@100% TPR), and the ablation study on the set-based MI process is in Appendix C.2.
@@
Figure 12: Performance evaluation under different image blending ratios (left) and noise perturbation budgets (right) (5 target samples per user). Zoom in to view the samples marked with different intensities.
C.1.2 Data marking with different parameters
There are two parameters: the blending ratio $m$ and the noise perturbation budget $\sigma$. Our main experiments use $m=0.7$ and $\sigma=8/255$, and we now study how different parameters may affect MembershipTracker’s performance. The results are in Fig. 12, where we consider a more challenging case with 5 target samples per user, because the performance difference between different marking intensities is less distinctive when there are more target samples by each user (e.g., 15).
@@ C.1.2 Data marking with different parameters
C.1.3 Alternative data marking methods
@@
(2) Alternative noise injection. We compare injecting perlin noise vs. uniform random noise. We find that perlin noise yields better results (details omitted), because perlin noise is a stronger adversarial noise [17], and thus more capable of inducing the model to memorize the marked samples. Other alternatives such as using other noises like gabor noise [17] may also be explored in future studies.
Figure 13: MembershipTracker maintains similarly high performance when using different types of outlier features for data marking, such as samples with random color stripes, samples from an OOD dataset (TinyImageNet and CelebA).
C.2 Set-based MI Verification
@@
Table 9: Comparing the instance-based and the proposed set-based MI process. The latter consistently outperforms the former method by leveraging the average loss across the target samples by each user to perform MI auditing.
@@
Appendix D Comparison with Wenger et al. [68]
Understanding the problematic evaluation in Wenger et al [68]. In their work, a testing feature should be detected as being used for training, if it causes higher probability shift than the non-target feature - this procedure should be the same regardless of whether the testing feature has actually been used for training or not (i.e., true or false positive). Unfortunately, this is not the case in [68].
To compute false positives (Line 2), however, Wenger et al. [68] simply invert the order of Line 1, which essentially compares the probability shift between the non-target feature $T_{out}^{\prime}$ and the true target feature ($T_{in}$). This would report no false positives (because $T_{out}^{\prime}$ has a lower probability shift than $T_{in}$), but it is not the correct way to compute false positive.
Instead, a fair evaluation should: (1) sample a new non-target feature ($T_{out}$); and (2) compare its probability shift relative to another non-target feature ($T_{out}^{\prime}$) [^3] In fact, the order of $T_{out}$ and $T_{out}^{\prime}$ in Line 3 of Alg. 2 can be interchangeable, as they are both random features that are not used for training. We use the given order only to illustrate how Wenger et al. [68] could overlook a false positive in their original detection.. This is illustrated in Line 3 of Alg. 2 (which is basically the same as Line 1 for detecting true positive). In such a rectified procedure, the technique would have a false positive, because $T_{out}$ has a higher probability shift than $T_{out}^{\prime}$.
@@
Algorithm 2 Illustrating the problematic evaluation in measuring false positive by [68], and how to fix it.
Validation. To evaluate how this affects their results, we use the implementation from the authors [68]. We use CIFAR100 dataset with 50 target classes and 50 non-target features for comparison; and otherwise follow the the same parameters they used in their experiment [68]. We first validate that we are able to obtain similar results to their reported performance, with 99.63% TPR and 0% FPR [^4] We identified a bug in their code (an issue related to pass-by-reference in Python that results in the code using a different set of auxiliary data for marking the target and non-target features), and we fixed it in our evaluation. We have also notified the authors about this bug in their official code repository using an anonymous account..
@@
Wenger et al. [68] do mention that comparing two sets of non-target features is equivalent to random guessing, and we also observe the same. However, this “indistinguishablity” applies only when attempting to distinguish two set of features; if we consider two random individual features, it is likely that one feature will cause a higher/lower probability shift than the other. This is the exact issue with Wenger et al.’s method, which performs each test on a feature-to-feature level (Line 3 in Alg. 2). Thus, it still permits a large number of FPs.
How to reduce (and control) FPR? One way to reduce the FPR in [68] is to employ a tighter significance level $\lambda$, which is used to detect whether the probability shift caused by the target feature is statistically higher than the non-target feature, and it defaults to be 0.1 in [68]. However, this will also degrade the TPR, e.g., when we change to use a $\lambda$ of 1e-8, the FPR is reduced to 1.9% [^5] We also checked that larger $\lambda$ like 1e-10 cannot further reduce FPR., and the TPR is already degraded to $86.4\%$ [^6] Another option to reduce FPR is to adjust the proportion of positive tests ($\delta$) for the tool to return a positive outcome (details omitted): we increased $\delta$ from 0.6 (their default setting) to 1.0, and still observed a similar trade-off.. In addition, while using a tighter significance level can reduce the FPR, it still cannot control the FPR within a specific low regime (e.g., 0.01%).
To fulfill this requirement, a principled way is to adapt a similar setup established in existing MI studies [70, 8, 6]. This procedure is sketched in Line 4 to Line 10 in Alg. 2. However, based on our results in the previous paragraph, it is likely that the TPR by their technique would further decrease when the TPR is controlled at a lower FPR regime (e.g., 0.01%).
Further comparison with [68]. We further compare Wenger et al. with MembershipTracker in our setting, where we evaluate how effective is their approach, if adapted for MI-based data provenance.
Specifically, we consider 100 target users and each user marks 25 samples (0.1% of the training set). As in Wenger et al. [68], we use ImageNet images as the outlier features to be blended into the CIFAR100 samples, and MembershipTracker uses random color stripes as before. [68] have a configurable parameter to control the intensity of the outlier feature when overlaid to the target data - we thus compare MembershipTracker with their approach under varying marking intensities.
@@
Figure 14: Comparing the technique by Wenger et al. (under varying data marking intensities) and MembershipTracker in terms of their MI effectiveness for tracing data provenance.
@@
As shown in Fig. 14, compared with the baseline approach without any data marking, the method by Wenger et al. can indeed increase the MI success. For instance, the green line in Fig. 14 (marked using $m=0.6$, which is equivalent to the default marking intensity in [68]) increases the AUC from 0.7927 to 0.9253. However, this still has very limited performance in the low FPR regime. Increasing the marking intensity in their approach can further improve the MI success, but it: 1) comes at the cost of higher image distortion; and 2) still has lower auditing performance than MembershipTracker.
@@
Our earlier discussion in Section 4.3.2 explains that injecting outlier features alone is not enough, as the original features hinder the model’s memorization on the outlier features. To overcome this, we propose a novel approach to inject a small amount of perlin noise, which works similar to adversarial perturbation and can significantly enhance the model’s memorization on the marked samples (as evaluated earlier in Appendix C.1.1) while preserving high image quality (the fourth column in Fig. 15).
Combined with the proposed set-based MI process (another key component in our work), MembershipTracker is able to achieve perfect MI success (the blue line in Fig. 14) and outperforms the method by Wenger et al.
Figure 15: Samples marked with different intensities by the method of Wenger et al. vs. the samples marked by MembershipTracker. Our approach incurs significantly lower distortion to images than Wenger et al., and still outperforms their approach with much higher MI success.
@@
Summary. While both Wenger et al. and our work aim at detecting the unauthorized use of personal data in training DL models, their original technique greatly underestimates the detection false positives, and thus still struggles with achieving high TPR and low FPR.
@@
We also realign their technique for MI-based provenance. We find MembershipTracker outperforms their approach by: 1) incurring lower distortion to the data; and 2) achieving significantly higher auditing performance. This renders MembershipTracker a more practical data provenance tool.
Appendix E Comparison with Tramer et al. [65]
@@
This section compare MembershipTracker with another related work (Tramer et al). Their work proposes a technique to improve the MI success based on data poisoning. The main idea is to inject mislabeled samples into the model’s training set, which can transform the target samples into outliers and amplify their influence to the model’s decision, thereby making the target samples easier to be de-identified.
@@
Tramer et al. is the closest work to ours that can improve the MI success and without requiring the expensive shadow-model calibration. However, it assumes the users can mislabel their data, which can be challenging in the real-world scenarios where the users do not have control over the data labels [21, 68, 19]. In comparison, MembershipTracker makes no such assumption and it only requires the users to mark their data with some small and targeted changes, which is a more realistic setting.

### v2

@@ 6. Evaluation on ImageNet Training
MembershipTracker can still maintain 0% FPR@100% TPR, when the marking ratio is reduced from 0.01% to 0.005%. However, when the marking ratio is further reduced to 0.0025%, MembershipTracker experiences lower auditing performance - this trend is similar to that in our ablation study in Appendix C.1.1 and thus we defer the detailed discussion to Appendix C.1.1 (due to space constraints). With that said, even under such a challenging case, MembershipTracker still achieves reasonably good performance as shown in the blue line of Fig. 11. To the best of our knowledge, MembershipTracker is the first technique that can scale to support reliable data auditing in large-scale training setting.
Figure 11. MembershipTracker under full-sized ImageNet training with different portions of samples for data marking. Aggressively reducing the portion of samples marked by each user can lead to reduced performance, though MembershipTracker can still maintain reasonably high auditing performance even when each user can only mark merely 0.0025% of the training data (the blue line).
7. Discussion
In this section, we first discuss how our assumption on the access to a set of non-member samples can be relaxed, and then discuss the limitations of MembershipTracker.
7.1. (Relaxed) Assumption on Non-member Set
Do non-member samples need to be definitely outside the training set? Recall that MembershipTracker follows a common assumption (Carlini et al., 2022; Bertran et al., 2024; Wenger et al., 2024; Ye et al., 2022) that users have access to a set of non-member samples (for calibrating the MI process under low FPR regime). However, in practice, it can be challenging for users to determine whether these non-member samples are definitely not used for training the target model. For instance, an auditing user may collect some random facial images from the Internet as non-member samples, but it is possible that these samples already have been used for training the target model.
To understand how this may affect MembershipTracker’s performance, we study a hypothetical scenario where all the non-member samples (used for data auditing) have been used for training the target model. Even under such a challenging case, we find that MembershipTracker can still maintain reliable auditing performance with 0% FPR@100% TPR. We explain the reason below.
Recall that MembershipTracker performs data marking on non-members using random outlier features. These random features are highly unlikely to be used for training (e.g., among $10^{16}$ possible random color stripe features), and thus they significantly increase the prediction loss of non-member samples, e.g., reducing their accuracy from $>$99% (this is high because these non-member samples are used for training) to $<$30% after data marking. Consequently, the marked non-member samples still exhibit much higher prediction loss than marked member samples, from which MembershipTracker still maintains 0% FPR@100% TPR for reliable data auditing.
Overall, we find that with MembershipTracker, users only need to collect some data from the underlying population, and these data can be any data; they do not necessarily have to be outside the training set of the target model. This demonstrates the broader practicality of MembershipTracker.
7.2. Limitations
Next, we discuss the three main limitations by MembershipTracker.
@@ 7.2. Limitations
MembershipTracker works by marking a small number of target samples to be strongly memorized by the model, and detecting whether the model memorizes these samples. However, a malicious developer may seek to jeopardize MembershipTracker’s performance, by using only a subset of data from each user for training, so that not all target samples are memorized by the model.
Table 5. Evaluation on the partial inclusion of user data in model training. With lesser data (50 samples per user) included for training, the average loss on the entire set of samples by each target user increases. This in turn (1) increases the FPR incurred by MembershipTracker, and (2) leads to larger accuracy drop due to lesser data for training.
% of user samples FPR@100% TPR Accuracy drop
used for model training
@@
90% (45/50) 0.00 0.12
70% (35/50) 0.00 4.83
50% (25/50) 0.34 12.34
40% (20/50) 0.84 14.76
30% (15/50) 2.42 21.37
To understand how this affects MembershipTracker, we perform an experiment where we assume the model owner collects 50 samples from each user (0.2% of the training set), and he/she includes only a subset of samples from each user for training. Table 5 shows the results, where we consider five target users (we observe a similar trend when considering the support for more target users). The FPR incurred by MembershipTracker grows as the number of marked samples used for training decreases. This is because those samples that are not used for training would yield higher prediction loss, which in turn increases the average loss on the entire set of marked samples, and affects MembershipTracker. On the other hand, while using a subset of user data can degrade MembershipTracker’s performance, it also results in lesser data for model training and causes an accuracy drop (third column in Table 5).
Overall, we find that MembershipTracker can still maintain $<0.5\%$ FPR@100% TPR, even when only 50% of the user data are used for training. In this case, the model already suffers from a 12% accuracy drop. Using lesser amounts of user data for training can further degrade MembershipTracker’s auditing performance, but it will also result in a larger accuracy drop (undesirable).
@@
Despite our best effort in keeping the targeted changes from severely distorting the marked data, the artifacts created by MembershipTracker are still visible under close inspection (more visual examples are shown in Appendix A). However, performing manual inspection to filter out the specially-marked samples by MembershipTracker can still be a challenge, particularly under the large-scale training setting (e.g., with millions of images). With that said, creating more subtle targeted changes still remains an important venue for future investigation.
@@
Finally, since MembershipTracker targets the model’s propensity in memorizing data, in principle, any effective measures that can mitigate the model’s memorization on training data would be able to degrade the performance by MembershipTracker. Indeed, our evaluation in Section 5.2 comprehensively considers multiple measures applied to model, input, and output levels; we find that though some approaches such as differential privacy can be configured to degrade the auditing success achieved by MembershipTracker, they also result in utility loss to the model (undesirable).
@@
8. Conclusion and Future Work
This work presents MembershipTracker, a practical data provenance tool that can support ordinary users to audit whether their specific data are used to train deep learning models without their permission. MembershipTracker consists of a lightweight data marking process that can mark the target data with small and targeted changes, and a high-power set-based membership inference process that can reliably trace the provenance of the target data. By merely marking a small fraction of samples (0.005$\sim$0.1of the dataset), MembershipTracker effectively empowers the users to reliably audit the usage of their data (average 0% FPR@100% TPR), and it is also scalable to the large-scale (ImageNet) training setting. If their data are found to be misused, the users can take legal action or request the “right to be forgotten” in accordance with privacy regulations such as GDPR (gdp, [n. d.]) (e.g., based on techniques on machine unlearning (Bourtoule et al., 2021)).
Finally, we discuss future work directions. First, while MembershipTracker retrofits the conventional membership inference (attack) to enable responsible AI, it also faces the risk of being misused by the malicious party. Notably, the data marked with MembershipTracker would now become easier to be de-identified by any party with access to them. However, MembershipTracker should be applied only when the users do not want their data to be used for training DL models, in which membership privacy may not be of the users’ concern. By contrast, if the users are willingly sharing their data for model training, they should refrain from using MembershipTracker to avoid unnecessary membership exposure. Future studies can explore potential usage cases where MembershipTracker is applicable and protecting the membership privacy of the marked samples is necessary, and consider the further adaptation of MembershipTracker to cater to such a need.
Second, future work can explore the potential of extending MembershipTracker to other domains like generative models. Large generative models are often trained on massive amount of data collected from the Internet (Rombach et al., 2022; Ramesh et al., 2022), and are the frequent targets where unauthorized use of personal/copyrighted contents arises (Dixit, 2023; Vincent, 2023). future work can investigate whether the large generative models would also be prone to memorize the samples marked with MembershipTracker under the billion-scale training environment.
@@ Appendix A Model Training Details
For CIFAR10 and CIFAR100, we train a WideResNet-28-4 model on 25,000 samples for 100 epochs, with a learning rate of 0.1 (decayed by 5 at epoch 60, 80, 90), a weight decay of 5e-4 and momentum of 0.9. For TinyImageNet and ArtBench, we fine-tune an ImageNet-pretrained ResNet-18 model on 25,000 samples. For CelebA, we first follow (cel, [n. d.]) to create a 307-class subset from the dataset, and then similarly fine-tune an ImageNet-pretrained model on 2,000 samples. For fine-tuning, we use a small fine-tuning rate of 0.01 with 30 epochs and a momentum of 0.9. In all datasets except CelebA (where we find that fine-tuning on the entire set yields better accuracy), we use 20% of the training data as the validation set. We also use common data augmentations such as random cropping and horizontal flip.
@@
Table 6. Evaluation on different models (parenthesized numbers show accuracy difference compared with the original models).
@@
Table 7. Evaluation on different training-set sizes.
@@ Appendix C Ablation Study
MembershipTracker consists of a data marking and set-based MI process, and we conduct a detailed ablation study into these two in Appendix C.1 and C.2, respectively.
C.1. Ablation of the Two-step Data Marking
@@
Table 8. Ablation study of MembershipTracker’s two-step data marking process: 1ga1g: Image blending with OOD feature; 1gb1g: Perlin noise injection. Each column reports the performance under a specific number of target samples contributed by the users (5$\sim$15 samples by each).
@@
(2) From left to the right: MembershipTracker’s auditing performance degrades as the number of samples contributed by the users reduces. However, even in the challenging case where each user marks only 5 samples, the full MembershipTracker still incurs only 0.64% FPR@100% TPR - this highlights the importance of data marking in MembershipTracker.
@@
C.1.1. Investigating the two-step data marking process
@@
We consider a total of 100 target users, and each contributes different number of samples (5 to 15) to the training set (25,000 instances). Table 8 shows the results, and there are two main findings.
@@
Meanwhile, the performance difference between different approaches in Table 8 shrinks as the number of target samples contributed by the users increases. Thus, there is a low FPR even when auditing the original target samples (without data marking), e.g., 0.34% FPR@100% TPR on the setting with 15 target samples per user. This is in fact due to the effectiveness of the proposed set-based MI process [^5] The ablation study on MembershipTracker with and without the set-based MI process is in Appendix C.2. However, this is not enough and using the two-step data marking still achieves better performance. The performance advantage of performing data marking is even more profound under the more challenging cases with limited user samples (third and fourth column in Table 8).
@@
Figure 13. Performance evaluation under different image blending ratios (left) and noise perturbation budgets (right) (5 target samples per user). Zoom in to view the samples marked with different intensities.
C.1.2. Data marking with different parameters
There are two parameters: the blending ratio $m$ and noise perturbation budget $\sigma$. Our main experiments use $m=0.7$ and $\sigma=8/255$, and we study how different parameters may affect MembershipTracker’s performance. The results are in Fig. 13, where we consider a more challenging case with 5 target samples per user, because the performance difference between different marking intensities is less distinctive when there are more target samples by each user (e.g., 15).
@@ C.1.2. Data marking with different parameters
C.1.3. Alternative data marking methods
@@
(2) Alternative noise injection. We compare injecting perlin noise vs. uniform random noise. We find perlin noise yields better results (details omitted), because perlin noise is a stronger adversarial noise (Co et al., 2019), and thus more capable of inducing the model to memorize the marked samples. Other alternatives such as using other noises like gabor noise (Co et al., 2019) may also be explored in future studies.
Figure 14. MembershipTracker maintains similarly high performance when using different types of outlier features for data marking, such as samples with random color stripes, samples from an OOD dataset (TinyImageNet and CelebA).
C.2. Ablation of the Set-based MI Verification
@@
Table 9. Comparing the instance-based and the proposed set-based MI process. The latter consistently outperforms the former method by leveraging the average loss across the target samples by each user to perform MI auditing.
@@
Algorithm 2 Illustrating the problematic evaluation in measuring false positive by (Wenger et al., 2024), and how to fix it.
Appendix D Comparison with Wenger et al. (Wenger et al., 2024)
Analyzing the problematic evaluation in Wenger et al (Wenger et al., 2024). In their work, a testing feature should be detected as being used for training, if it causes higher probability shift than the non-target feature - this procedure should be the same regardless of whether the testing feature has actually been used for training or not (i.e., true or false positive). Unfortunately, this is not the case in (Wenger et al., 2024).
To compute false positives (Line 2), however, Wenger et al. (Wenger et al., 2024) simply invert the order of Line 1, which essentially compares the probability shift between the non-target feature $T_{out}^{\prime}$ and the true target feature ($T_{in}$). This would report no false positives (because $T_{out}^{\prime}$ has a lower probability shift than $T_{in}$), but it is not the correct way to compute false positive.
Instead, a fair evaluation should: (1) sample a new non-target feature ($T_{out}$); and (2) compare its probability shift relative to another non-target feature ($T_{out}^{\prime}$) [^6] In fact, the order of $T_{out}$ and $T_{out}^{\prime}$ in Line 3 of Alg. 2 can be interchangeable, as they are both random features that are not used for training. We use the given order only to illustrate how Wenger et al. (Wenger et al., 2024) could overlook a false positive in their original detection.. This is illustrated in Line 3 of Alg. 2 (which is basically the same as Line 1 for detecting true positive). In such a rectified procedure, the technique would have a false positive, because $T_{out}$ has a higher probability shift than $T_{out}^{\prime}$.
Validation. To evaluate how this affects their results, we use the original implementation from (Wenger et al., 2024). We use CIFAR100 dataset with 50 target classes and 50 non-target features for comparison; and otherwise follow the the same parameters they used in their experiment (Wenger et al., 2024). We first validate that we can obtain similar results to their reported performance, with 99.63% TPR and 0% FPR [^7] We identified a bug in their code (an issue related to pass-by-reference in Python that results in the code using a different set of auxiliary data for marking the target and non-target features), and we fixed it in our evaluation. We have also notified the authors about this bug in their official code repository using an anonymous account..
@@
Wenger et al. (Wenger et al., 2024) do mention that comparing two sets of non-target features is equivalent to random guessing, and we also observe the same. However, this “indistinguishablity” applies only when attempting to distinguish two set of features; if we consider two random individual features, it is likely that one feature will cause a higher/lower probability shift than the other. This is the exact issue with Wenger et al.’s method, which performs each test on a feature-to-feature level (Line 3 in Alg. 2). Thus, it still permits a large number of FPs.
How to reduce (and control) FPR? One way to reduce the FPR in (Wenger et al., 2024) is to employ a tighter significance level $\lambda$, which is used to detect whether the probability shift caused by the target feature is statistically higher than the non-target feature, and it defaults to be 0.1 in (Wenger et al., 2024). However, this will also degrade the TPR, e.g., when we change to use a $\lambda$ of 1e-8, the FPR is reduced to 1.9% [^8] We also checked that larger $\lambda$ like 1e-10 cannot further reduce FPR., and the TPR is already degraded to $86.4\%$ [^9] Another option to reduce FPR is to adjust the proportion of positive tests ($\delta$) for the tool to return a positive outcome (details omitted): we increased $\delta$ from 0.6 (their default setting) to 1.0, and still observed a similar trade-off.. In addition, while using a tighter significance level can reduce the FPR, it still cannot control the FPR within a specific low regime (e.g., 0.01%).
To fulfill this requirement, a principled way is to adapt a similar setup established in existing MI studies (Ye et al., 2022; Carlini et al., 2022; Bertran et al., 2024). This procedure is sketched in Line 4 to Line 10 in Alg. 2. However, based on our results in the previous paragraph, it is likely that the TPR by their technique would further decrease when the TPR is controlled at a lower FPR regime (e.g., 0.01%).
Further comparison with (Wenger et al., 2024). We further compare Wenger et al. with MembershipTracker in our setting, where we evaluate how effective is their approach, if adapted for MI-based data provenance.
Specifically, we consider 100 target users and each user marks 25 samples (0.1% of the training set). As in Wenger et al. (Wenger et al., 2024), we use ImageNet images as the outlier features to be blended into the CIFAR100 samples, and MembershipTracker uses random color stripes as before. (Wenger et al., 2024) have a configurable parameter to control the intensity of the outlier feature when overlaid to the target data - we thus compare MembershipTracker with their approach under varying marking intensities.
@@
Figure 15. Comparing the technique by Wenger et al. (under varying data marking intensities) and MembershipTracker in terms of their MI effectiveness for data auditing.
@@
Figure 16. Samples marked with different intensities by the method of Wenger et al. vs. the samples marked by MembershipTracker. Our approach incurs significantly lower distortion to images than Wenger et al., and still outperforms their approach with much higher MI success.
As shown in Fig. 15, compared with the baseline approach without any data marking, the method by Wenger et al. can indeed increase the MI success. For instance, the green line in Fig. 15 (marked using $m=0.6$, which is equivalent to the default marking intensity in (Wenger et al., 2024)) increases the AUC from 0.7927 to 0.9253. However, this still has very limited performance in the low FPR regime. Increasing the marking intensity in their approach can further improve the MI success, but it: 1) comes at the cost of higher image distortion; and 2) still has lower auditing performance than MembershipTracker.
@@
Our earlier discussion in Section 4.3.2 explains that injecting outlier features alone is not enough, as the original features hinder the model’s memorization on the outlier features. To overcome this, we propose a novel approach to inject a small amount of perlin noise, which works similar to adversarial perturbation and can significantly enhance the model’s memorization on the marked samples (as evaluated earlier in Appendix C.1.1) while preserving high image quality (the fourth column in Fig. 16). Combined with the proposed set-based MI process (another key component in our work), MembershipTracker is able to achieve perfect MI success (the blue line in Fig. 15) and outperforms the method by Wenger et al.
@@
Summary. While both Wenger et al. and our work aim at detecting the unauthorized data use in training DL models, their original technique greatly underestimates the detection false positives, and thus still struggles with achieving high TPR and low FPR.
@@
We also realign their technique for MI-based provenance. We find MembershipTracker outperforms their approach by: 1) incurring lower distortion to the data; and 2) achieving significantly higher auditing performance. This renders MembershipTracker a more practical data auditing tool.
Appendix E Comparison with Tramer et al. (Tramèr et al., 2022)
@@
This section compares MembershipTracker with another related work (Tramer et al). Their work proposes a technique to improve the MI success based on data poisoning. The main idea is to inject mislabeled samples into the model’s training set, which can transform the target samples into outliers and amplify their influence to the model’s decision, thereby making the target samples easier to be de-identified.
@@
Tramer et al. is the closest work to ours that can improve the MI success and without requiring the expensive shadow-model calibration. However, it assumes the users can mislabel their data, which can be challenging in the real-world scenarios where the users do not have control over the data labels (Dzieza, 2023; Wenger et al., 2024; Deng et al., 2009). In comparison, MembershipTracker makes no such assumption and it only requires the users to mark their data with some small and targeted changes, which is a more realistic setting.